A conic (E) has equation y2+2y+8x+9=0

a) State the nature of (E)

b) Determine its characteristic elements

c) Sketch (E) indicating all the elements.

d) Find the length of the lactus.



1
Expert's answer
2021-09-27T12:31:13-0400

a)


"E: y^2+2y+8x+9=0"

"y^2+2y+1+8x+8=0"

A conic (E) is a parabola. Vertex form:


"x=-\dfrac{1}{8}(y+1)^2-1"

b)

"p=-2, k=-1, h=-1"

Vertex: "(h,k)=(\u22121,\u22121)"

Focus: "(h+p,k)=(-3,-1)"

Eccentricity: "1"

Directrix: "x=h-p, x=1"

Latus rectum: "x=\u22123"

The length of the latus rectum: "8"

Axis of symmetry: "y=\u22121"

Focal parameter: "1-(-3)=4"

x-intercept: "y=0, x=-\dfrac{1}{8}(0+1)^2-1=-\dfrac{9}{8}"

 "(\u22129\/8,0)"

No y-intercepts.


(c)



d) Latus rectum: "x=\u22123"

Find ends of latus rectum


"-3=-\dfrac{1}{8}(y+1)^2-1"

"y+1=\pm4"

"(-3, -5), (-3, 3)"

The length of the latus rectum


"3-(-5)=8"